### v1

# TOP2024: an overview of experimental results

Abstract
The $17^{\text{th}}$ edition of the international workshop on top quark physics featured a diverse set of outstanding results. This note is an attempt to summarize the workshop from the experimental perspective and suggest ways forward for the future investigations. As it has not been possible to touch upon every single physics analysis in this note, interested reader are referred to individual proceedings for details.
Copyright attribution to authors.
This work is a submission to SciPost Phys. Proc.
License information to appear upon publication.
Publication information to appear upon publication.
Received Date
Accepted Date
Published Date
1 Introduction
Since its discovery in 1995 [1, 2], the heaviest known particle, the top quark, has been a focus of studies in the experimental and theoretical particle physics. The large mass of top quark results in a very short life time as well as a Yukawa coupling close to one. The particle is therefore an excellent candidate to study the bare quark properties, to understand the electroweak symmetry breaking, and to connect with new phenomena. In proton-proton (pp) collisions at the CERN LHC, top quarks are abundantly produced in pair ($\text{t}\bar{\text{t}}$) through strong interactions, and to lesser extent, single via electroweak processes (single-top). The production in both cases can be accompanied by extra particles, including jets, vector bosons, the Higgs boson, and even additional top quarks. Currently, most of the production modes are observed and/or precisely measured. From precise measurements, the top quark mass and properties are determined while additional parameters of the standard model of particle physics (SM) are extracted. At the same time, searches for new phenomena with top quarks together with the interpretation of top quark measurements have put strong constraints on theories beyond the standard model (BSM).
The latest measurements and searches in top quark physics at the LHC rely on about two decades of scientific research, acquired knowledge, and innovation. The state-of-the-art analysis techniques and particle identification, theory advancements, and the large data sample provided by the LHC, have enabled the ATLAS and CMS experiments to perform high precision measurements, observe extremely rare processes, and strongly rule out various BSM possibilities. Figure 1 shows the inclusive cross section measurements using pp collisions collected by the CMS experiment  [^1] Similar information is available in Ref. [3], performed by the ATLAS collaboration., covering few orders of magnitude for processes involving top quark(s) [4]. The majority of measurements are limited by systematic uncertainties. It must also be noted that associated to almost every results in Fig. 1, there are (multi)differential measurements, extraction of SM parameters, and interpretations in the context of BSM theories.
Figure 1: Summary of production cross section measurements involving top quarks by the CMS experiment [4]. Similar information is available in Ref. [3], performed by the ATLAS collaboration.
Considering different top quark production modes, the large cross section of $\text{t}\bar{\text{t}}$, $\sigma_{\text{t}\bar{\text{t}}}$, has made it a unique target for the additional measurements and interpretations. It was the first to measure at the start of the LHC Run 3 [5, 6], already with a high precision. In addition, CMS reported a measurements at $\sqrt{s}=5.02\,\text{TeV}$ [7], to compare with the corresponding ATLAS result [8] and to be used as a reference for top quark studies in heavy ion collisions at the same center-of-mass energy, $E_{\text{c.o.m}}$. The ATLAS collaboration presented a $\sigma_{\text{t}\bar{\text{t}}}$ measurements in proton-lead (pPb) collisions at $\sqrt{s}=8.16\,\text{TeV}$ [9] to compare with an earlier result by CMS [10]. New results from both experiments call for further discussions and comparisons of analysis details and simulation settings, which can be particularly interesting when heavy ions are involved.
The available knowledge and expertise paves the way for future investigations in the field of top quark physics. By the end of 2025, the full LHC data at $\sqrt{s}=13.6\,\text{TeV}$ will be available for physics analyses. The High-Luminosity LHC is expected to start in , providing the largest-ever amount of data. At the same time, experiments will undergo a substantial upgrade to sustain the immense luminosity and to fully exploit the wealth of the data. This is a unique opportunity to explore uncharted territories in top quark physics, drawing on the existing and developing knowledge.
In addition to a summary of the experimental results presented in TOP2024, this note provides a few suggestions to expand the research in the coming years. As a disclaimer, the article presents a personal view and does not cover all possibilities.
2 Machine learning for top quark measurements
[…]
Another example of the ML usage is to preserve the precision of top quark measurements where the limited size of Monte-Carlo (MC) samples, especially those with a varied parameter to assess the associated systematic uncertainty, is a limiting factor. Using ML techniques, the effect of the varied parameter can be estimated in the phase space of study with no need for large MC production [13]. This approach is crucial, especially for the high-luminosity LHC, where MC production can hardly cope with the amount of collected data for precise measurements. Other areas where ML has an impact on top quark physics include the reconstruction of the top quark event hypothesis, particularly in the presence of neutrinos and extra particles [14], and estimation of backgrounds [15].
3 Opportunities with $\text{t}\bar{\text{t}}$ differential measurements
Differential distributions of $\sigma_{\text{t}\bar{\text{t}}}$ have been measured since the early times of the LHC data taking and provided valuable inputs to the upcoming measurements and to the theory community. The well-known observed slope in the differential measurement of top quark transverse momentum $p_{\text{T}}$ with respect to theory predictions at the time triggered a chain of studies over different data-taking eras. Multidifferential measurements were performed to possibly localize the effect where at the same time, higher order calculations could explain (part of) the disagreements. The acquired knowledge from multidifferential measurements went beyond understanding the top quark $p_{\text{T}}$ distributions and used to extract SM parameters such as $\alpha_{\text{S}}$ and top quark mass, and to constrain the gluon density function of protons  [^3] Refer to this presentation for a comprehensive overview.. The state-of-the-art ML methods have made it possible to perform simultaneous differential measurements of many more observables with a higher granularity in other physics processes [16, 17]. It is desired to use the same techniques for differential $\sigma_{\text{t}\bar{\text{t}}}$ measurements.
Possibility for a $\text{t}\bar{\text{t}}$ bound state (aka toponium)
Heavy pseudoscalars decaying to $\text{t}\bar{\text{t}}$ can create a peak-dip structure in the $m_{\text{t}\bar{\text{t}}}$ distributions, if exist. In a search for such particles, using $\text{t}\bar{\text{t}}$ angular and spin correlation observables, CMS has observed an excess with a significance of more than $5\sigma$ at $m_{\text{t}\bar{\text{t}}}$ values close to the $\text{t}\bar{\text{t}}$ production threshold [18]. The excess is compatible with the presence of a toponium signal, using the available calculations in which the color-octet state is not accounted for. Figure 2 shows the observed and expected $m_{\text{t}\bar{\text{t}}}$ distributions in bins of spin correlation observable. The SM distributions with an injected topnium signal are fitted to and compared with the data in the lower panel. The pseudoscalar hypothesis is slightly favored by the data for the observed excess. A similar measurement by ATLAS together with more realistic theory predictions and proposals for complementary measurements seem necessary to better identify the nature of this signal.
Figure 2: Observed and expected $m_{\text{t}\bar{\text{t}}}$ distribution in bins of spin correlation observables shown for the dilepton (e and/or $\mu$) final state summed over lepton flavors and analysis eras. In the first panel, the data points with statistical error bars and predicted perturbative QCD backgrounds (colored histograms) are compared before the fit to the data, and the corresponding prefit uncertainty is shown with a gray band. In the second panel, the ratio is shown with the best fit normalization applied considering toponium contribution in the fit. There, the gray band shows the postfit uncertainty, and the respective signal is overlaid with its best fit model parameters. Adapted from Ref. [18].
Quantum entanglement outside the light cone
Using the $\Delta\phi$ between the two leptons from W bosons in $\text{t}\bar{\text{t}}$ decays, ATLAS and CMS experiments observed quantum entanglement between the two top quarks near the $\text{t}\bar{\text{t}}$ production threshold [19, 20]. The entanglement signal remained statistically significant in the presence of possible $\text{t}\bar{\text{t}}$ bound state. In another work [21], the CMS experiment exploited the full $\text{t}\bar{\text{t}}$ spin density matrix and observed entangled top quarks that in part of the phase space with high $m_{\text{t}\bar{\text{t}}}$ and low top quark scattering angles. This is particularly important because in this regions top quarks are expected to be largely outside each other’s light cone and the information cannot be transmitted with hidden variables  [^4] Possibilities to experimentally probe quantum entanglement in single-top were presented here..
4 Associated top quark productions
While the latest $\text{t}\bar{\text{t}}$$\text{b}\bar{\text{b}}$ measurements by ATLAS and CMS [22, 23] show similar trends in comparison with theory predictions, comparing of the results is not straight forward because of different final states and definitions of the fiducial phase space (see Fig. 3 for the ATLAS result). In addition to harmonized selections, common MC samples with similar generators and scale choices would be extremely useful in this context. The CMS $\text{t}\bar{\text{t}}$$\text{c}\bar{\text{c}}$ measurement has to be updated with the high-performance c-taggers, and in view of the latest ATLAS result [24]. The LHC searches for $\text{t}\bar{\text{t}}$$\text{t}\bar{\text{t}}$ have been concluded with the observation of the process in Run 2, with the possibility to combination the results of the two experiments with higher precisions in Run 3. The off-shell $\text{t}\bar{\text{t}}$$\text{t}\bar{\text{t}}$ production has shown constraining power on the Higgs boson width and top-Higgs Yukawa coupling $\kappa_{\text{t}}$ [25]. Given the sensitivity of differential $m_{\text{t}\bar{\text{t}}}$ and $\Delta y_{\text{t}\bar{\text{t}}}$ measurements to $\kappa_{\text{t}}$ [26], there is more room for top quark physics to contribute to the Higgs sector of the SM.
Figure 3: Measured fiducial cross sections for compared with the central values of NLO $\text{t}\bar{\text{t}}$$\text{b}\bar{\text{b}}$ predictions from different event generators and settings [22]. The right hand panel shows the ratios of the MC predictions to the data. A similar trend is observed in the CMS measurement [23] performed in lepton+jets final states and compared with different event generators and settings.
The vector boson associated production of top quark involves $\text{t}\bar{\text{t}}$, tVq, and tWV, with V here being a Z boson or a photon. Precise $\text{t}\bar{\text{t}}\mathrm{\gamma}$ measurements exist where experiments need to harmonize the definition of the visible phase space, and the origin of the photon. The current $\text{t}\mathrm{\gamma}\text{q}$ observation has to be followed by differential measurements, preferably simultaneously with $\text{t}\bar{\text{t}}\mathrm{\gamma}$. To date, there are no dedicated $\text{t}\text{W}\mathrm{\gamma}$ measurements at the LHC while an evidence has been found for the tWZ production [27]. In view of future tWV measurements, the treatment of quantum interference with $\text{t}\bar{\text{t}}$ has to advance in simulation, ideally accounting for non-, single- and double-resonant processes at NLO QCD in a single MC sample. A simultaneous differential measurement of $\text{t}\bar{\text{t}}Z$ and tZq has been recently reported by CMS [28]. It provides a consistent treatment of the two processes, which is particularly important for constraining top quark electroweak interactions in the context of the effective field theory (EFT).
5 BSM searches and interpretations
The top quark is a key player of BSM searches including dark matter, vector-like quarks, and lepton flavor and baryon number violations  [^5] Details were presented in [1, 2, 3]. . In addition, there are huge efforts in both theory and experimental communities to constrain, with maximum information and minimum assumptions, the new couplings in the EFT Lagrangian using top quark measurements. Recently, CMS has released a global EFT fit based on likelihood functions from a wide range of measurements [29]. This can be confronted with the biggest global fit from the theory side [30] that constrains 50 Wilson coefficients using 455 data points. The experimental fit is going to expand with more event signatures, include latest modeling of the processes, and employ ML techniques. An additional step can be the inclusion of searches that are interpreted in the EFT framework [31]. The theory fit can also improve by using more of the available differential measurements, especially those of associated top quark productions.
6 Conclusion
The top quark physics at the LHC advances relying on the outstanding work from theory and experiment. Exchanges across the communities are vital for the continuation of the research. On the experimental side, the succesful experience of common Monte-Carlo $\text{t}\bar{\text{t}}$ samples between ATLAS and CMS can be extended to other top quark processes to facilitate the comparison of definitions and the results.
Acknowledgements
The author thanks the organizers of the workshop for the invitation and for the excellent coordination.
Funding information
The participation of the author is financially supported by the Ministry for Research, Science, and Technology of Iran, and by Isfahan University of Technology.